OpenStudy (anonymous):

A grocer wants to mix two kinds of nuts. One kind sells for $0.95 per pound, and the other sells for $2.60 per piund. He wants us to mix a total of 16 pounds and sell it for $2.10 per pound. How many pounds of each kind should he use in the new mix? (round off answers to the nearers hundredth.)

OpenStudy (amistre64):

see if i can recall my formula :) b = t(z-x) at y ----- (y-x) b at .95 = 16(2.10 - 2.60) -------------- (.95 - 2.60)

OpenStudy (amistre64):

that gets me; rounded to (4.85) pounds of the cheap stuff.. but let me dbl chk

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

yep; thats works :) 4.85 of the cheap stuff 11.15 of the good stuff

OpenStudy (anonymous):

\[.95x+2.6(16-x)=16\times 2.1\] solve and get what sensei said